OpenStudy (kayleruthless):

What is the vertex of the graph of f(x) = |x + 5| – 6? (–6, –5) (–6, 5) (–5, –6) (5, –6)

OpenStudy (sshayer):

x+5=0 x=-5 when |x+5|=0,y=f(x)=-6